# Graph Triple Attention Network: A Decoupled Perspective

### v1

@@ Abstract.
Graph Transformers (GTs) have recently achieved significant success in the graph domain by effectively capturing both long-range dependencies and graph inductive biases. However, these methods face two primary challenges: (1) multi-view chaos, which results from coupling multi-view information (positional, structural, attribute), thereby impeding flexible usage and the interpretability of the propagation process. (2) local-global chaos, which arises from coupling local message passing with global attention, leading to issues of overfitting and over-globalizing. To address these challenges, we propose a high-level decoupled perspective of GTs, breaking them down into three components and two interaction levels: positional attention, structural attention, and attribute attention, alongside local and global interaction. Based on this decoupled perspective, we design a decoupled graph triple attention network named DeGTA, which separately computes multi-view attentions and adaptively integrates multi-view local and global information. This approach offers three key advantages: enhanced interpretability, flexible design, and adaptive integration of local and global information. Through extensive experiments, DeGTA achieves state-of-the-art performance across various datasets and tasks, including node classification and graph classification. Comprehensive ablation studies demonstrate that decoupling is essential for improving performance and enhancing interpretability. Our code is available at: https://github.com/wangxiaotang0906/DeGTA.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Recently, Graph Transformers (GT) (Dwivedi and Bresson, 2020; Hussain et al., 2022; Rampášek et al., 2023) have shown great potential in handling graph-structured data such as social networks (Qiu et al., 2018), drug discovery (Stokes et al., 2020), and traffic networks (Derrow-Pinion et al., 2021). The success of GTs is built upon their ability to leverage inductive bias and handle long-range dependencies simultaneously. For example, Graphormer (Ying et al., 2021) utilizes global attention to capture long-range dependencies and uses structural biases to influence attention scores, thereby introducing inductive bias. GraphGPS (Rampášek et al., 2023) employs GNNs before global attention to enhance graph topology information. Besides, these methods will compensate positional encodings and structural encodings into node features to enhance the performance. However, existing GTs mainly face two challenges:
@@
2) local-global chaos: Existing GTs couple local message passing and global attention, which complicates the adjustment of importance weights between local and global information. This uniformity potentially precipitates issues of overfitting and over-globalizing (Xing et al., 2024). In some graphs, local information may be crucial, while in others, global information may dominate. Coupling these sources of information makes it challenging to adapt the weights appropriately and lacks interpretability regarding which information contributes to success.
@@
  • We identify the main challenges of GTs and propose a decoupled perspective that provides a framework for designing novel decoupled models.
@@ 2.1. Graph Attention Networks & GTs
The above models conduct message passing based on graph topology, e.g., MPNN-based (Gilmer et al., 2017; Kipf and Welling, 2017; Zhu et al., 2022, 2024a), which can be considered a local interaction between nodes. It’s well known that MPNN-based GNNs face challenges of over-smoothing (Alon and Yahav, 2021), oversquashing (Topping et al., 2022), and limited expressive power limitations (Oono and Suzuki, 2020). To extend the receptive field, Graph Transformers (GTs) (Dwivedi and Bresson, 2020; Hussain et al., 2022) have emerged, utilizing global attention combined with positional and structural information to simultaneously capture long-range dependencies and graph inductive bias. For instance, Graphormer (Ying et al., 2021) uses centrality encoding to enhance structural information and spatial matrix as positional biases, SAN (Kreuzer et al., 2021) uses a learnable Laplacian PE as input to a branch Transformer layer, and GRIT (Ma et al., 2023) uses a random walk encoding to integrate positional information. GraphGPS (Rampášek et al., 2023) explicitly integrate various other types of MPNN modules into their architectures. Other works such as NAGphormer (Chen et al., 2023) and SAT (Chen et al., 2022) use subgraph to introduce neighborhood inductive biases to node features. However, GTs face two significant problems: (1) local-global chaos, and (2) multi-view chaos. These issues make it challenging for GTs to adapt weights appropriately, lack interpretability regarding which information contributes to success, and impede flexible usage and the interpretability of the propagation process.
@@ 3. Revisiting Graph Attention Mechanism: A Decoupled Perspective
3.1. Decoupled Perspective
3.1.1. Decoupled Perspective of Multi-View Attentions
(iii) Attribute Encoding (AE) aim to transform raw node attributes, such as text or images, into numeric vectors. Recently, with the significant success of LLMs in language understanding, text-attributed graphs (TAGs) have become a prominent topic in the graph domain. For TAGs, LLMs can be employed to enhance attribute encodings (He et al., 2023; Zhu et al., 2024b). Thus we can select AE from original features (Fey and Lenssen, 2019) or LLM-enhanced features (Zhu et al., 2024b; Wang et al., 2024).
Previous methods, like GraphGPS (Rampášek et al., 2023) and Graphormer (Ying et al., 2021), couple these information types to calculate attention, leading to multi-view chaos, which impedes flexible usage and the interpretability of the propagation process. The clear definitions can guide us in checking if the information is adequate and help address the multi-view chaos of attention in graphs by separately computing structural, positional, and attribute attention through different encodings of node pairs.
3.1.2. Decoupled Perspective of Message Interaction
@@ 4.3. Decoupling Message Interaction (RQ2)
(9) $\hat{H}_{i}^{global}=\sum_{j\in\mathcal{K}_{i}}\hat{z}_{ij}H_{j},\;\text{where}\;\hat{z}_{i,j}=\frac{e^{z_{i,j}}}{\sum_{k\in\mathcal{K}_{i}}e^{z_{i,k}}}\$
@@ 4.4. Discussion
4.4.2. Analysis for Expressivity and Over-smoothing.
@@
Additionally, we provide both empirical and theoretical evidence demonstrating that DeGTA offers stronger resistance to over-smoothing compared to GT. Details are in Appendix D.
@@ 5. Experiments
In this section, we present a comprehensive empirical investigation. Specifically, we aim to address the following research questions: RQ1: How does the proposed DeGTA perform in node classification tasks, including both homophilic and heterophilic datasets? RQ2: How does DeGTA perform in graph-level tasks? RQ3: How effectively does DeGTA capture long-range dependencies? RQ4: How does the decoupling of multi-view and local-global influence performance? Additionally, we provide a parameter study on $K$ in Appendix A.2, and the case study that demonstrates our enhanced expressive power and interpretability in Figure 6.
@@ 5.3.2. Experiments on Long-range Dependencies Datasets (RQ3)
5.4. Ablation Study (RQ4)
@@ Appendix A More Experiment details
Table 7. Details of node classification datasets.
Dataset #Nodes #Edges #Features #Classes Split(%)
Pubmed 19,717 44,324 500 3 0.3/2.5/5.0
Citeseer 3,327 4,552 3,703 6 5.2/18/37
Cora 2,708 5,728 1,433 7 3.6/15/30
Arxiv 169,343 1,166,243 128 40 53.7/17.6/28.7
Texas 183 279 1,703 5 48/32/20
Cornell 183 277 1,703 5 48/32/20
Wisconsin 251 450 1,703 5 48/32/20
Actor 7,600 26,659 932 5 48/32/20
AMiner-CS 593,486 6,217,004 100 18 20/30/50
Amazon2M 2,449,029 61,859,140 100 47 20/30/50
@@
ZINC 12,000 23 25 regression - Mean Abs. Error 83.3/8.3/8.3
MINST 70,000 71 565 classif. 10 Accuracy 78.6/7.1/14.2
CIFAR10 60,000 118 941 classif. 10 Accuracy 75/8.3/16.7
Peptides-func 15,535 151 307 classif. 10 Avg. Precision 70/15/15
Peptides-struct 15,535 151 307 regression 11 Mean Abs. Error 70/15/15
@@ A.1. Details of Datasets
The Pubmed, Citeseer, and Cora datasets are citation networks (Yang et al., 2016). Each node represents a research article and two nodes are adjacent if there is a citation between two articles. The node attributes are the bag-of-words features, and the node label is the category of the research domain of the article.
The ogbn-arxiv (Hu et al., 2020) dataset constitutes a directed graph typifying the citation network amid all Computer Science (CS) arXiv manuscripts as cataloged by the Microsoft Academic Graph (MAG). Each vertex within this graph is an arXiv document, while each directed edge signifies a citation from one document to another. Accompanying each paper is a 128-dimensional feature vector, derived from the mean of the embeddings corresponding to words in the document’s title and abstract sections.
The Texas, Cornell, and Wisconsin datasets are extracted from the WebKB dataset (Pei et al., 2020). Each node represents a webpage and two nodes are adjacent if there is a hyperlink between the two webpages. The node attributes are the bag-of-words features, and the node label is the category of the webpage.
The actor dataset is the actor-only induced subgraph of a film-director-actor-writer network obtained from Wikipedia webpages (Pei et al., 2020). Each node represents an actor and two nodes are adjacent if the two corresponding actors appear on the same Wikipedia webpage. The node features are derived from the keywords on the Wikipedia webpage of the corresponding actor, and the node label is determined by the words on the webpage.
AMiner-CS (Feng et al., 2020) are citation networks in which nodes represent papers and edges represent citations. Amazon2M (Chiang et al., 2019) are co-purchase networks, where nodes indicate goods and edges indicate that the two connected goods are frequently bought together. The splits of large-scale datasets are followed the settings from (Feng et al., 2022).
ZINC (Dwivedi et al., 2020) consists of 12K molecular graphs from the ZINC database of commercially available chemical compounds. These molecular graphs consist of 9 to 37 nodes that each represents a heavy atom (28 possible atom types) and each edge represents a bond (3 possible types). The task is to regress constrained solubility (logP) of the molecule.
MNIST and CIFAR10 (Dwivedi et al., 2020) are derived from like-named image classification datasets by constructing an 8 nearest-neighbor graph of SLIC superpixels for each image.
Peptides-func and Peptides-struct (Dwivedi et al., 2022b) are composed of atomic graphs of peptides retrieved from SATPdb. In Peptides-func the prediction is multi-label graph classification into 10 nonexclusive peptide functional classes. While for Peptides-struct the task is graph regression of 11 3D structural properties of the peptides.
@@ A.2. Details of hyperparameter
Experimental results are reported on the hyperparameter settings below, where we choose the settings that achieve the highest performance on the validation set. We choose hyperparameter grids that do not necessarily give optimal performance, but hopefully cover enough regimes so that each model is reasonably evaluated on each dataset.
  • lr $\in\{5e-2,1e-2,5e-3,1e-3,5e-4\}$
  • $K$$\in\{2,3,4,6,8,12\}$
  • pe_dim ($d_{p}$) $\in\{2,4,6,8,12,16\}$
  • se_dim ($d_{s}$) $\in\{2,4,6,8,12,16\}$
  • ae_dim ($d$) $\in\{32,64,128,256,512\}$
  • pe&se_hidden_dim ($d^{\prime}$) $\in\{2,4,6,8,12,16\}$
  • ae_hidden_dim ($d^{\prime\prime}$) $\in\{32,64,128,256,512\}$
  • dropout $\in\{0,0.1,0.2,0.3,0.5,0.8\}$
  • weight_decay $\in\{1e-2,5e-3,1e-3,5e-4,1e-4\}$
  • activation $\in\{\text{elu},\text{relu},\text{leakyrelu}\}$
  • layer_num $\in\{1,2,3,4,5,6,7,8\}$
@@ Appendix B Details of Initialization Strategies of PE/SE
In this section, we provide a detailed description of the encoding initialization strategies discussed in Section 1.
@@
0:  Normalized adjacency matrix $\tilde{A}$; Positional receptive field $K$
@@ Appendix C Expressive power of DeGTA
1-Weisfeiler-Leman (1-WL) Test and MPNNs. The 1-WL test is a node-coloring algorithm within the hierarchy of WL heuristics used for graph isomorphism. It operates by iteratively updating node colors based on their 1-hop local neighborhoods until no further changes occur in the node colors. The limitations of MPNNs in distinguishing non-isomorphic graphs were rigorously analyzed in the work of (Xu et al., 2019). This analysis highlights the established equivalence between MPNNs and the 1-WL isomorphism test, as detailed by (Morris et al., 2019). As a result, MPNNs may perform poorly on graphs that exhibit multiple symmetries within their original structure, including node and edge isomorphisms.
@@
Appendix D Over-smoothing analysis of DeGTA
In this section, we provide both experimental and theoretical analyses of DeGTA’s ability to mitigate the over-smoothing problem. Firstly, we introduce a quantitative metric for smoothness. We use the average Euclidean distance of the node attributes as a measure of smoothness and calculate this metric at each layer as follows:
$\begin{split}sm(G)&=\frac{1}{n}\sum_{i\in V}\frac{1}{n-1}\sum_{j\in(V-i)}D(x_{i},x_{j})\\ &=\frac{1}{n}\sum_{i\in V}\frac{1}{n-1}\sum_{j\in V}\frac{1}{2}\left|\left|\frac{x_{i}}{\parallel x_{i}\parallel}-\frac{x_{j}}{\parallel x_{j}\parallel}\right|\right|\\ &=\frac{1}{2n(n-1)}\sum_{i}\sum_{j}\left|\left|\frac{x_{i}}{\parallel x_{i}\parallel}-\frac{x_{j}}{\parallel x_{j}\parallel}\right|\right|\end{split}$
[…]
Experimental results. We conducted experiments on three datasets, evaluating accuracy and smoothness of DeGTA and GT as the number of layers increases. As shown in Figure 7, the experimental results demonstrate that DeGTA significantly enhances model performance and alleviates over-smoothing at each layer compared to GT, showing that the hard sampling strategy of DeGTA effectively captures the important long-range dependencies while avoiding the harm associated with the high degree of freedom in global information propagation of GT.
Theoretical analysis. In the following, we prove theoretically that DeGTA is further away from over-smoothing than GT when have equal attention to long-range dependencies.
First, we give the overall node update formula for DeGTA, where $N(i)$ is the set of neighboring nodes and $K(i)$ is the set of sampled nodes. For simplicity, we omit the details of the model here and frozen the adaptive integration of local-global. Ultimately, our node update can be written as ${A_{N}}_{i}X+{A_{S}}_{i}X$, where $A_{N}$ only pays attention to neighboring nodes and $A_{S}$ only pays attention to sampled nodes.
$\hat{x}_{i}=\sum_{j\in N(i)}a_{ij}x_{j}+\sum_{j\in K(i)}a_{ij}x_{j}$
There is a normalization method so that each row of A is equal to 1, which leads to that when GT and DeGTA have the same attention to important long-range dependencies, the nodes of DeGTA have relatively higher attention to themselves as well as to their neighboring nodes than GT, focusing on important long-range dependencies while better preserving their own information and avoiding the phenomenon of over-smoothing.
$sm(\hat{G})_{DeGTA}$ $=\sum_{i,j}\left|\left|\frac{\hat{x}_{i}}{||\hat{x}_{i}||}-\frac{\hat{x}_{j}}{||\hat{x}_{j}||}\right|\right|_{DeGTA}$
$=\sum_{i,j}\big{|}\big{|}\hat{x}_{i}-\hat{x}_{j}\big{|}\big{|}_{DeGTA}$
$=\sum_{i,j}\sum_{x}\big{|}\big{|}\hat{x}_{ix}-\hat{x}_{jx}\big{|}\big{|}_{DeGTA}$
$=\sum_{i,j}\sum_{x}\bigg{|}\bigg{|}\sum_{y\in N(x)}x_{iy}{A_{N}}_{yx}+\sum_{y\in K(x)}x_{iy}{A_{S}}_{yx}-$
$\sum_{y\in N(x)}x_{jy}{A_{N}}_{yx}-\sum_{y\in K(x)}x_{jy}{A_{S}}_{yx}\bigg{|}\bigg{|}$
$=\sum_{i,j}\sum_{x}\bigg{|}\bigg{|}\sum_{y\in N(x)}(x_{iy}-x_{jy}){A_{N}}_{yx}+\sum_{y\in K(x)}(x_{iy}-$
$x_{jy}){A_{S}}_{yx}\bigg{|}\bigg{|}$
$=\sum_{i,j}\sum_{x}\sum_{y\in N(x)}\big{|}\big{|}(x_{iy}-x_{jy})\big{|}\big{|}{A_{N}}_{yx}+\sum_{i,j}\sum_{x}\sum_{y\in K(x)}$
$\big{|}\big{|}(x_{iy}-x_{jy})\big{|}\big{|}{A_{S}}_{yx}$
$\geq\sum_{i,j}\sum_{x}\sum_{y\in N(x)\cup K(x)}\big{|}\big{|}(x_{iy}-x_{jy})\big{|}\big{|}({A_{N}}_{yx}+{A_{S}}_{yx})$
$\geq\sum_{i,j}\sum_{x}\sum_{y\in V}\big{|}\big{|}(x_{iy}-x_{jy})\big{|}\big{|}A_{yx}$
$=\sum_{i,j}\sum_{x}\left|\left|\sum_{y\in V}x_{iy}A_{yx}-\sum_{y\in V}x_{jy}A_{yx}\right|\right|$
$=\sum_{i,j}\big{|}\big{|}\hat{x}_{i}-\hat{x}_{j}\big{|}\big{|}_{GT}$
$=sm(\hat{G})_{GT}$
The inequality holds when the nodes of DeGTA and GT have the same attention to long-range dependencies in $K(i)$, the $sm$ metric of DeGTA is larger than that of GT, proving that our DeGTA captures the same long-range dependencies while being further away from over-smoothing than GT.
Figure 7. The results of the smoothness of DeGTA compared with GT of each layer. DeGTA demonstrates higher performance while maintaining further away from over-smoothing at each layer.

### v2

@@ Abstract.
Graph Transformers (GTs) have recently achieved significant success in the graph domain by effectively capturing both long-range dependencies and graph inductive biases. However, these methods face two primary challenges: (1) multi-view chaos, which results from coupling multi-view information (positional, structural, attribute), thereby impeding flexible usage and the interpretability of the propagation process. (2) local-global chaos, which arises from coupling local message passing with global attention, leading to issues of overfitting and over-globalizing. To address these challenges, we propose a high-level decoupled perspective of GTs, breaking them down into three components and two interaction levels: positional attention, structural attention, and attribute attention, alongside local and global interaction. Based on this decoupled perspective, we design a decoupled graph triple attention network named DeGTA, which separately computes multi-view attentions and adaptively integrates multi-view local and global information. This approach offers three key advantages: enhanced interpretability, flexible design, and adaptive integration of local and global information. Through extensive experiments, DeGTA achieves state-of-the-art performance across various datasets and tasks, including node classification and graph classification. Comprehensive ablation studies demonstrate that decoupling is essential for improving performance and enhancing interpretability.
Graph Neural Network, Graph Transformer, Graph Representation Learning
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709223isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Artificial intelligenceccs: Information systems Data mining
@@ 1. Introduction
Recently, Graph Transformers (GT) (Dwivedi and Bresson, 2020; Hussain et al., 2022; Rampášek et al., 2023) have shown great potential in handling graph-structured data such as social networks (Qiu et al., 2018), drug discovery (Stokes et al., 2020), and traffic networks (Derrow-Pinion et al., 2021). The success of GTs is built upon their ability to leverage inductive bias and handle long-range dependencies simultaneously. For example, Graphormer (Ying et al., 2021) utilizes global attention to capture long-range dependencies and uses structural biases to influence attention scores, thereby introducing inductive bias. GraphGPS (Rampášek et al., 2023) employs GNNs before global attention to enhance graph topology information. Besides, these methods will compensate positional encodings and structural encodings into node features to enhance the performance. However, existing GTs mainly face two challenges while integrating diverse information:
@@
2) local-global chaos: Existing GTs couple local message passing and global attention, which complicates the adjustment of importance weights between local and global information. This uniformity potentially precipitates issues of overfitting and over-globalizing (Xing et al., 2024). In some graphs, local information may be crucial, while in others, global information may dominate. Coupling these sources of information makes it challenging to adapt the weights appropriately and lacks interpretability of which information contributes to success.
@@
  • We propose a decoupled perspective for attention in graphs. From this perspective we identify two main challenges of information coupling in GTs and provide a framework for designing novel decoupled models.
@@ 2.1. Graph Attention Networks & GTs
The above models conduct message passing based on graph topology, e.g., MPNN-based (Gilmer et al., 2017; Kipf and Welling, 2017; Zhu et al., 2022; Wang et al., 2024), which can be considered a local interaction between nodes. It’s well known that MPNN-based GNNs face challenges of over-smoothing (Alon and Yahav, 2021), oversquashing (Topping et al., 2022), and limited expressive power limitations (Oono and Suzuki, 2020). To extend the receptive field, Graph Transformers (GTs) (Dwivedi and Bresson, 2020; Hussain et al., 2022) have emerged, utilizing global attention combined with positional and structural information to simultaneously capture long-range dependencies and graph inductive bias. For instance, Graphormer (Ying et al., 2021) uses centrality encoding to enhance structural information and spatial matrix as positional biases, SAN (Kreuzer et al., 2021) uses a learnable Laplacian PE as input to a branch Transformer layer, and GRIT (Ma et al., 2023) uses a random walk encoding to integrate positional information. GraphGPS (Rampášek et al., 2023) explicitly integrate various other types of MPNN modules into their architectures. Other works such as NAGphormer (Chen et al., 2023) and SAT (Chen et al., 2022) use subgraph to introduce neighborhood inductive biases to node features. However, GTs face two significant problems while integrating different categories of information: (1) local-global chaos, and (2) multi-view chaos. These issues make it challenging for GTs to adapt weights appropriately, lack interpretability regarding which information contributes to success, and impede flexible usage and the interpretability of the propagation process.
@@ 3. Revisiting Graph Attention Mechanism: A Decoupled Perspective
3.1. Decoupled Perspective
3.1.1. Decoupled Perspective of Multi-View Attentions
(iii) Attribute Encoding (AE) aim to transform raw node attributes, such as text or images, into numeric vectors. Recently, with the significant success of LLMs in language understanding, text-attributed graphs (TAGs) have become a prominent topic in the graph domain. For TAGs, LLMs can be employed to enhance attribute encodings (He et al., 2023; Zhu et al., 2024b). Thus we can select AE from original features (Fey and Lenssen, 2019) or LLM-enhanced features (Zhu et al., 2024b, a).
Previous methods like GraphGPS (Rampášek et al., 2023) and Graphormer (Ying et al., 2021), couple these information types to calculate attention, leading to multi-view chaos, which impedes flexible usage and the interpretability of the propagation process. The clear definitions can guide us in checking if the information is adequate and help address the multi-view chaos of attention in graphs by separately computing structural, positional, and attribute attention through different encodings of node pairs.
3.1.2. Decoupled Perspective of Message Interaction
@@ 4.3. Decoupling Message Interaction (RQ2)
(9) $\hat{H}_{i}^{\text{global}}=\sum_{j\in\mathcal{K}_{i}}\hat{z}_{ij}H_{j},\;\text{where}\;\hat{z}_{i,j}=\frac{e^{z_{i,j}}}{\sum_{k\in\mathcal{K}_{i}}e^{z_{i,k}}}\$
@@ 4.4. Discussion
4.4.2. Analysis for Expressivity.
@@ 5. Experiments
In this section, we present a comprehensive empirical investigation. Specifically, we aim to address the following research questions: RQ1: How does the proposed DeGTA perform in node classification tasks, including both homophilic and heterophilic datasets? RQ2: How does DeGTA perform in graph-level tasks? RQ3: How effectively does DeGTA capture long-range dependencies? RQ4: How does DeGTA enhances the interpretability of message aggregation? RQ5: How does the decoupling of multi-view and local-global influence performance? Additionally, we provide a parameter study on $K$ and the sensitivity analysis of selection of PE/SE in Appendix A.2. Moreover, a case study that demonstrates our enhanced expressive power and interpretability is in Figure 6.
@@ 5.3.2. Experiments on Long-range Dependencies Datasets (RQ3)
5.4. The Enhanced Interpretability (RQ4)
The enhancement of multi-view interpretability provided by DeGTA is twofold.
  • Node level. As shown in the case study (Figure 6), visualizing the intermediate results of multi-view attention separately enables us to identify which information is more similar across node pairs.
  • Graph level. As shown in Table 6, analyzing the training outcomes of different attention weights allows us to discern which types of information are most critical for the specific dataset.
As shown in Table 6, structural and attribute weights dominate in heterophilic datasets, while positional information is minimal. This indicates that global message, attribute, and structural information are vital in heterophilic graphs, whereas local message and positional information are more beneficial in homophilic graphs.
Table 6. The adaptive weights of different datasets after training.
Dataset Homophily Positional weight Structural weight Attribute weight
Pubmed 0.66 0.5432 0.0763 0.3805
Citeseer 0.63 0.3479 0.1022 0.5499
Cora 0.77 0.5759 0.0827 0.3414
Arxiv 0.42 0.2893 0.1374 0.5733
Texas 0.00 0.0729 0.1406 0.7865
Cornell 0.02 0.0263 0.1663 0.8074
Wisconsin 0.05 0.1331 0.2347 0.6342
Actor 0.01 0.0281 0.2156 0.7563
5.5. Ablation Study (RQ5)
@@ 6. Conclusion
Acknowledgments
This work was supported by the National Key Research and Development Plan of China (2023YFB4502305), and Ant Group through Ant Research Intern Program.
@@ Appendix A More Experiment details
ZINC (Dwivedi et al., 2020) 12,000 23 25 regression - Mean Abs. Error 83.3/8.3/8.3
MINST (Dwivedi et al., 2020) 70,000 71 565 classif. 10 Accuracy 78.6/7.1/14.2
CIFAR10 (Dwivedi et al., 2020) 60,000 118 941 classif. 10 Accuracy 75/8.3/16.7
Peptides-func (Dwivedi et al., 2022b) 15,535 151 307 classif. 10 Avg. Precision 70/15/15
Peptides-struct (Dwivedi et al., 2022b) 15,535 151 307 regression 11 Mean Abs. Error 70/15/15
@@ A.2. Details of hyperparameter
Experimental results are reported on the hyperparameter settings in our code link, where we choose the settings that achieve the highest performance on the validation set. We choose hyperparameter grids that do not necessarily give optimal performance, but hopefully cover enough regimes so that each model is reasonably evaluated on each dataset.
@@
Sensitivity analysis of the selection of PE/SE. DeGTA is robust and insensitive to the selection of PE/SE and the multi-view encoder. We implemented a simple combination of Jaccard/RWSE (well-established in previous works), utilizing MLP as the encoder for various inputs, yet achieved substantial performance.
Table 10 demonstrating that arbitrary combinations of encodings can surpass the performance of coupled approaches within the DeGTA framework. Thus we not selectively choose from existing methods to find a combination that perform well, but offer a guiding decoupled framework which is robust to all settings.
Table 10. Ablation experiment of different PE/SE in Arxiv. DeGTA is insensitive to the choice of encoding for multi-view information, but sensitive to decoupling.
PE SE DeGTA (decoupling) coupling (PE+SE) coupling (AE+PE)
RWPE RWSE 73.31 71.69 71.38
RWPE DSE 72.97 70.99 70.94
Jaccard RWSE 73.26 72.13 72.02
Jaccard DSE 73.12 72.08 70.82
LapPE RWSE 73.09 71.24 71.36
LapPE DSE 72.54 70.98 69.97
@@ Appendix B Details of Initialization Strategies of PE/SE
In this section, we provide a detailed description of the encoding initialization strategies discussed in Section 3.1.1.
@@
0:  Normalized adjacency matrix $\tilde{A}$; Receptive field $K$
@@ Appendix C Expressive power of DeGTA
1-Weisfeiler-Leman (1-WL) Test and MPNNs. The 1-WL test is a node-coloring algorithm within the hierarchy of WL heuristics used for graph isomorphism. The limitations of MPNNs in distinguishing non-isomorphic graphs were rigorously analyzed in the work of (Xu et al., 2019). This analysis highlights the established equivalence between MPNNs and the 1-WL isomorphism test, as detailed by (Morris et al., 2019). As a result, MPNNs may perform poorly on graphs that exhibit multiple symmetries within their original structure, including node and edge isomorphisms.